what is the 20% of $192.15 and how?
Answer by stanbon(75887)

20% = 20/100 = 0.20
----
what is the 20% of $192.15 and how?
-----
Equation:
0.20*192.15 = 0.20*1.9215 = $38.43
=======================
Cheers,
Stan H.
=======================